One number is 8 more than anothwr . Their sum is -12. Finf the numbers.
Answer by solve_for_x(190)

Let x represent the smaller of the two numbers.

The larger number is then represented by x + 8.

Since the sum of the two numbers is -12, we can write:

x + (x + 8) = -12

2x + 8 = -12

2x = -20

x = -10

The smaller number is -10.

The larger number is then -10 + 8 = -2

Solution: The numbers are -2 and -10.